OpenStudy (anonymous):

An arc of length 10 meters is formed by a central angle A on a circle of radius 4. The measure of A in degrees (to two decimal places) is _____. A. 216.76 B. 143.24 C. 2.50 D. 36.76

OpenStudy (jdoe0001):

bear in mind that \(\bf s = \cfrac{\theta r \pi}{180} \implies \theta = \cfrac{180 s}{r \pi}\)

OpenStudy (anonymous):

so 180s/12.56?

OpenStudy (anonymous):

Hello I am your teacher and I will be calling your mom :)